16 lb of Brand M Cinnamon was made by
combining 12 lb of Indonesian cinnamon
which costs $19/lb with 4 lb of Thai
cinnamon which costs $11/lb. Find the cost
per lb of the mixture.

Answer by richwmiller(17219)

12*19+4*11=16*x
228+44=16*x
272=16x
x=$17 lb